OpenStudy (anonymous):

Simplify: 12(5c - 3).

OpenStudy (anonymous):

\[12(5c) -12(3) = 60c-36\]

OpenStudy (anonymous):

60c-36

OpenStudy (aravindg):

60c-36

OpenStudy (anonymous):

remember \[a(b+c) =ab+ac\]

OpenStudy (anonymous):

60c-36